# Phase diagram of two-component mean-field Bose mixtures

### v1

@@ I Introduction
Bose mixtures received an explosion of interest in the years following the realization of Bose-Einstein condensation (BEC) in trapped ultracold atomic gases, and their intense exploration continues up to now [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Despite this, a number of key features concerning the phase diagram of such systems have been unveiled only very recently. For the most elementary case of interaction- and mass-balanced mixtures with simple repulsive interactions, these include the first-order character of the transition to the BEC states occurring in substantial portions of the phase diagram, and previously unexplored liquid-gas type phase transitions within the regions of the phase diagram occupied by the non-condensed phases. It seems plausible that density jumps associated with such transitions may be detectable in current cold-atom experiments, in particular considering that their exploration does not require cooling the system below the condensation temperature. Apart from the above-mentioned aspects of the phase diagram, which were neglected or overlooked in the early theoretical studies, there are other properties of Bose mixtures, which, up to date, remain hardly explored theoretically and are also likely to become of growing interest and relevance in near future experimental investigations. These concern in particular the case of negative (attractive) interspecies coupling and the limit, where the system is driven towards collapse [34, 35, 36, 37, 38] by tuning interactions. Another aspect of Bose mixtures that has not been explored much that has not been explored much relates to the effect of interaction and mass imbalance on the phase diagram structure as well as thermodynamic and excitation spectra properties.
@@ II Model and review of previous results
Here $i,j\in\{1,2\}$, the intraspecies couplings $a_{i,i}>0$ (hereafter denoted as $a_{i}$) are assumed positive, while the interspecies coupling $a_{1,2}=a_{2,1}$ (denoted hereafter as $a_{12}$) has no restriction on its sign. We will, in particular, analyze the instability occurring when $a_{12}$ is negative and becomes sufficiently large in absolute value. The single-particle states for each type of particles are labelled by the wavevector $\vec{k}$, $\hat{n}_{\vec{k},i}$ are the corresponding occupation number operators, $\hat{N}_{i}:=\sum_{\vec{k}}\hat{n}_{\vec{k},i}$ denote the total particle number operators for each of the species, and the dispersion takes the standard form $\epsilon_{\vec{k},i}=\hbar^{2}\vec{k}^{2}/(2m_{i})$. The system is subject to periodic boundary conditions, maintained at fixed temperature $k_{B}T=\beta^{-1}$ and enclosed in a three-dimensional cubic vessel of volume $V$. We will use the grand canonical ensemble, where the chemical potentials $\mu_{1}$, $\mu_{2}$ of the mixture constituents are fixed, and the grand-canonical free energy density $\omega(T,\mu_{1},\mu_{2})$ follows from the grand canonical partition function $\Xi(T,V,\mu_{1},\mu_{2})$ via
@@
The construction of the model represents a rather straightforward extension of the so-called imperfect Bose gas model [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51] to the case of binary mixtures. We note in particular that the mean-field interaction term $\sim N^{2}/V$ may be derived from a realistic two-body interaction potential via the so-called Kac scaling procedure [52], where the interaction is progressively scaled to be extremely weak and long-ranged at the same time. We also point out [32] that the structure of the self-consistency (saddle-point) equations deriving from the model of Eq. (1) is, up to redefinition of parameters, equivalent to the one arising from the Hartree-Fock treatment [19] of the weakly interacting Bose gas in the dilute limit. We therefore expect that the conclusions of the present study are not restricted to the setup involving extremely long-ranged interactions and may be pertinent to realistic situations.
@@
[…]
@@
[…]
@@
where the terms proportional to $1/V$ represent the condensates’ densities for $V\to\infty$. Analogously, the function $\Phi$ given by Eq. (5) can be recast as
@@
Notice that the last two terms in the above expression always vanish in the thermodynamic limit, in contrast to their counterparts in Eqs (9,10), which yield the condensate densities. We can use Eq. (9,10) to determine $n_{1}$ as a function of $n_{2}$ (or vice versa). Within the normal phase, we have:
@@
Importantly, for certain parameter ranges (see [32]), the system admits multiple solutions and identifying the physically realized equilibrium requires checking which of them corresponds to the lowest value of $\Phi$; the remaining solutions then represent metastable or unstable equilibria. The previous analysis of Ref. [32], largely by numerical means, explored the general structure of the phase diagram in the interaction- and mass-balanced case with repulsive interactions. The major findings of that study are summarized as follows. For $D>0$, the phase diagram hosts, in addition to the normal phase, a phase involving condensate of only type-1 particles (BEC1), only type-2 particles (BEC2) and condensates of both species (BEC12), where all the involved phase transitions are continuous. This situation is schematically depicted in Fig. 1 (the right panel). In contrast, for $D<0$, there is no BEC12 phase, and the transition between BEC1 and BEC2 is of first order and is accompanied by phase separation; condensation of single species of particles may occur either via a first order or continuous transition depending on system parameters. In addition, for sufficiently large, positive interspecies couplings $a_{12}$, there occurs a line of liquid-gas-type phase transitions within the normal phase; for a schematic illustration, see Fig. 1 (left panel). This picture was primarily obtained through numerical analysis of Eq. (5,9,10) for situations with neither mass nor interaction imbalance and repulsive interparticle forces. We presently systematize, generalize, and extend that analysis to parameter ranges not addressed before, also providing a much more thorough analytical understanding. We begin by discussing the possibility of attractive couplings.
@@ III Attractive interspecies interactions
We first observe that the one-component counterpart of the system given by Eq. (1) is well-defined only for non-negative values of the interaction coupling. In regard to the presently studied two-component mixture, we note that the structure of the saddle-point equations [Eq. (9,10)] does not impose immediately obvious constraints on the values of the interaction couplings $\{a_{1},a_{2},a_{12}\}$. To identify restrictions on the values of these parameters stemming from the system stability, we reexamine the structure of the integrals of Eq. (3,4). We first point out [45, 46, 32] that the result is independent of the arbitrary parameters $\alpha_{1}$ and $\alpha_{2}$ occurring in the integration limits. These parameters are taken to be negative to guarantee that the function $\Phi(t_{1},t_{2})$ can (at least for repulsive interactions) be analytically continued into the region of the complex plane encompassing the integration contour. To ensure that the Bose function $g_{5/2}(x)$ and the logarithmic function $\ln(x)$ are well-defined, both $|e^{\beta t_{1}}|$ and $|e^{\beta(\frac{a_{12}}{a_{1}}t_{1}+t_{2})}|$ must be less than one. This condition is satisfied provided that the real parts of $t_{1}$ and $(\frac{a_{12}}{a_{1}}t_{1}+t_{2})$ are negative. When $D>0$, it is straightforward to see that this can be guaranteed by appropriate choice of $\alpha_{1}$ and $\alpha_{2}$ (irrespective of the sign of $a_{12}$). On the other hand, when $D<0$ and $a_{12}<0$, it is impossible to fulfil the above conditions. To see this, we express $t_{1}$ as the sum of its real and imaginary parts, $t_{1}=a+ib$. Fulfillment of the inequality $|e^{\beta t_{1}}|<1$ requires that $a<0$. However, the condition $|e^{\beta(\frac{a_{12}}{a_{1}}t_{1}+t_{2})}|<1$ requires $a$ to be positive, as $a_{12}<0$ and $t_{2}$ is integrated over the entire real axis. This leads to a contradiction, rendering it impossible to satisfy both conditions simultaneously when $D<0$. In consequence, the integrals in Eq. (3, 4, 5) remain well defined also for attractive interspecies interactions but only as long as $D>0$. We also note that the intraspecies couplings $a_{1}$, $a_{2}$ must necessarily be positive.
@@
Figure 2: Projection of the phase diagram on the $(\mu_{1},\mu_{2})$ plane for a sequence of values of the interspecies coupling $a_{12}$ [$\beta a_{12}\lambda_{1}^{-3}=0.9$ (blue), $\beta a_{12}\lambda_{1}^{-3}=0.1$ (red), and $\beta a_{12}\lambda^{-3}_{1}=-0.7$ (grey)] for sufficiently weak interspecies coupling $a_{12}$ such that $D>0$. The BEC12 phase occupies the wedge-shaped region characterized by large values of $\mu_{1}$ and $\mu_{2}$ (corresponding also to high densities). The system hosts a quadruple point, where the BEC12, BEC1, BEC2 and the normal phase coexist. The BEC12 wedge opens upon decreasing $a_{12}$, its angle reaches the value $\phi=90$°  for $a_{12}=0$ and approaches $\phi=180$°  in the limit $a_{12}\to-\sqrt{a_{1}a_{2}}^{+}$ (where $D\to 0^{+}$ and the system loses its stability). The plot parameters are $\beta a_{1}\lambda_{1}^{-3}=\beta a_{2}\lambda_{1}^{-3}=\lambda_{1}/\lambda_{2}=1$. All the phase transitions are continuous (see Sec. IV).
For $D>0$ and relatively small and positive values of $a_{12}$, the BEC12 phase is stable in the acute wedge-shaped region characterized by large values of $\mu_{1}$ and $\mu_{2}$ (which also correspond to high densities) - see Fig. 2. The system hosts a quadruple point, where the BEC12 phase coexists with the BEC1, BEC2, as well as with the normal phase. All the phase transitions involved here are continuous (see Sec. IV). Upon reducing $a_{12}$, the position of the quadruple point becomes shifted to lower values of $(\mu_{1},\mu_{2})$, and the BEC12 wedge progressively opens. The wedge angle $\phi$ reaches the value 90°  for $a_{12}=0$ and is further increased as $a_{12}$ becomes negative. The limiting value $a_{12}\to-\sqrt{a_{1}a_{2}}^{+}$ (where $D\to 0^{+}$) corresponds to the fully open wedge (ϕitalic-$\phi\to$180°). This picture may be understood realizing that at the BEC2-BEC12 transition one has [see Eq. (9, 10)]:
@@ IV Tricritical points for condensation
In this section, we analyze the conditions for realizing condensation as a first-order transition. We demonstrate the impossibility of such a scenario for $D>0$ (a fact used in the analysis of the previous section). We also show that $D<0$ is a condition necessary, but not sufficient, for appearance of first-order condensation in the system’s phase diagram. This scenario will however always be realized for $T$ sufficiently low.
We focus on the transition line between the normal and BEC2 phases; the normal-BEC1 case may be addressed in a fully analogous manner. For sufficiently large negative values of $\mu_{1}$ ($-\beta\mu_{1}\gg 1$), the concentration of type-1 particles is negligibly low and properties of the mixture are equivalent to those of the one-component system. In particular, there is only one solution to the saddle-point equations; condensation of type-2 particles occurs via a continuous transition, and the corresponding phase boundary follows Eq. (14). It may, however, happen [32] that, as one moves along the line defined by Eq. (14), additional solutions to the saddle-point equation appear and Eq. (14) begins to represent an unstable equilibrium. This scenario requires an appearance of a tricritical point in the phase diagram ($P_{2,c}$ in Fig. 1), which is reflected by a change of the character of the extremum of $\Phi(n_{1}(n_{2}),n_{2})$ at the critical density. This, in turn, implies an existence of a point on the line given by Eq.(14) where the second derivative of $\Phi(n_{1}(n_{2}),n_{2})$ with respect to $n_{2}$ vanishes. Using Eq. (13), we view $\Phi(n_{1},n_{2})$, as a function of $n_{2}$ only and we introduce $\tilde{\Phi}(n_{2})\equiv\Phi(n_{1}(n_{2}),n_{2})$. The condition for vanishing of the second derivative of $\tilde{\Phi}(n_{2})$ then may be written as:
@@
We emphasize that the above discussion implies that the tricritical points may occur exclusively for negative values of the determinant $D$. However, a negative $D$ does not necessarily imply that Eqs. (20, 21) indeed represent a meaningful tricritical point on the phase diagram. The additional requirement is that the putative tricritical point of Eq. (20, 21) be located to the left (i.e. correspond to lower values of the chemical potentials) from the point of intersection of the putative second-order phase boundaries [given by Eq. (14) and an analogous equation for the normal-BEC1 transition]. Numerical analysis indicates that this condition becomes fulfilled for sufficiently low temperatures and/or $D$ sufficiently close to zero. This is easily proven in the balanced case, where $a_{1}=a_{2}=a$ and $\lambda_{1}=\lambda_{2}=\lambda$. In this case, the condition for Eq. (20, 21) to represent a physical tricritical point reduces to $\mu_{1,tri}<\mu_{2,tri}$, which may be recast as

### v2

@@ I Introduction
Bose mixtures received an explosion of interest in the years following the realization of Bose-Einstein condensation (BEC) in trapped ultracold atomic gases, and their intense exploration continues up to now [27, 20, 50, 65, 55, 5, 24, 62, 58, 3, 47, 6, 53, 12, 4, 23, 28, 11, 46, 21, 68, 69, 44, 13, 42, 54, 59, 14, 70, 2, 37, 40, 66, 7, 9, 8, 51, 52, 25, 56, 29, 39, 61, 30, 57, 63, 64, 43, 34, 35, 15]. Despite this, a number of key features concerning the phase diagram of such systems have been unveiled only very recently. For the most elementary case of interaction- and mass-balanced mixtures with simple repulsive interactions, these include the first-order character of the transition to the BEC states occurring in substantial portions of the phase diagram, and previously unexplored liquid-gas type phase transitions within the regions of the phase diagram occupied by the non-condensed phases. It seems plausible that density jumps associated with such transitions may be detectable in current cold-atom experiments, in particular considering that their exploration does not require cooling the system below the condensation temperature. Apart from the above-mentioned aspects of the phase diagram, which were neglected or overlooked in the early theoretical studies, there are other properties of Bose mixtures, which, up to date, remain hardly explored theoretically and are also likely to become of growing interest and relevance in near future experimental investigations. These concern in particular the case of negative (attractive) interspecies coupling and the limit, where the system is driven towards collapse [36, 60, 16, 1, 45] by tuning interactions. Another aspect of Bose mixtures that has not been explored much relates to the effect of interaction and mass imbalance on the phase diagram structure as well as thermodynamic and excitation spectra properties.
@@
Although many aspects of the phase diagram of homogeneous two-component Bose-Bose mixtures were studied before [27, 20, 50, 65, 55, 5, 24, 62, 58, 3, 47, 6, 53, 12, 4, 23, 28, 11, 46, 21, 68, 69, 44, 13, 42, 54, 59, 14, 70, 2, 37, 40, 66, 7, 9, 8, 51, 52, 25, 56, 29, 39, 61, 30, 57, 63, 64, 43, 34, 35, 15], the general picture emergent from available literature is scattered, largely incomplete, and occasionally misleading. In addition, our work offers general analytical insights, whereas the majority of earlier analyses relied on numerical results restricted to specific parameter sets.
@@ II Model and review of previous results
Here $i,j\in\{1,2\}$, the intraspecies couplings $a_{i,i}>0$ (hereafter denoted as $a_{i}$) are assumed positive, while the interspecies coupling $a_{1,2}=a_{2,1}$ (denoted hereafter as $a_{12}$) has no restriction on its sign. We will in particular, in Sec. III, analyze the instability occurring when $a_{12}$ is negative and becomes sufficiently large in absolute value. The single-particle states for each type of particles are labelled by the wavevector $\vec{k}$, $\hat{n}_{\vec{k},i}$ are the corresponding occupation number operators, $\hat{N}_{i}:=\sum_{\vec{k}}\hat{n}_{\vec{k},i}$ denote the total particle number operators for each of the species, and the dispersion takes the standard form $\epsilon_{\vec{k},i}=\hbar^{2}\vec{k}^{2}/(2m_{i})$. The system is subject to periodic boundary conditions, maintained at fixed temperature $k_{B}T=\beta^{-1}$ and enclosed in a three-dimensional cubic vessel of volume $V$. We will use the grand canonical ensemble, where the chemical potentials $\mu_{1}$, $\mu_{2}$ of the mixture constituents are fixed, and the grand-canonical free energy density $\omega(T,\mu_{1},\mu_{2})$ follows from the grand canonical partition function $\Xi(T,V,\mu_{1},\mu_{2})$ via
@@
The construction of the model represents a rather straightforward extension of the so-called imperfect Bose gas model [17, 10, 67, 41, 18, 71, 49, 48, 31, 33, 19, 32, 38, 22] to the case of binary mixtures. We note in particular that the mean-field interaction term $\sim N^{2}/V$ may be derived from a realistic two-body interaction potential via the so-called Kac scaling procedure [26], where the interaction is progressively scaled to be extremely weak and long-ranged at the same time. We also point out [34] that the structure of the self-consistency (saddle-point) equations deriving from the model of Eq. (1) is, up to redefinition of parameters, equivalent to the one arising from the Hartree-Fock treatment [68] of the weakly interacting Bose gas in the dilute limit. We therefore expect that the conclusions of the present study are not restricted to the setup involving extremely long-ranged interactions and may be pertinent to realistic situations.
@@
$\Xi(T,V,\mu_{1},\mu_{2})=-\frac{\beta V}{2\pi\sqrt{a_{1}^{\prime}a_{2}^{\prime}}}\int_{\alpha_{1}-i\infty}^{{\alpha_{1}+i\infty}}dt_{1}\int_{\alpha_{2}-i\infty}^{{\alpha_{2}+i\infty}}dt_{2}e^{-V\Phi(t_{1},t_{2})}\;,$ (3)
@@
$\Xi(T,V,\mu_{1},\mu_{2})=-\frac{i\beta V}{2\pi\sqrt{a^{\prime}_{1}|a^{\prime}_{2}}|}\int_{\alpha_{1}-i\infty}^{\alpha_{1}+i\infty}dt_{1}\int_{-\infty}^{\infty}dt_{2}e^{-V\Phi(t_{1},t_{2})}$ (4)
@@
where the terms proportional to $1/V$ represent the condensates’ densities for $V\to\infty$. The criterion for determining the presence or absence of condensates is based on the behaviour of these quantities in the thermodynamic limit as $V\to\infty$. Non-zero densities of particles in the one-particle ground states indicate the existence of condensates.
Analogously, the function $\Phi$ given by Eq. (5) can be recast as
@@
Notice that the last two terms in the above expression always vanish in the thermodynamic limit, in contrast to their counterparts in Eqs (9,10), which yield the condensate densities. The thermodynamic state is determined by solving the self-consistency equations [see Eq. (9) and Eq. (10)] and checking whether the one-particle ground states are macroscopically occupied. When more than one self-consistent solutions exist, the physically realized state corresponds to the dominant saddle point, which is the solution where $\Phi$ has the lowest value. We can use Eq. (9,10) to determine $n_{1}$ as a function of $n_{2}$ (or vice versa). Within the normal phase, we have:
@@
Importantly, for certain parameter ranges (see [34]), the system admits multiple solutions and identifying the physically realized equilibrium requires checking which of them corresponds to the lowest value of $\Phi$; the remaining solutions then represent metastable or unstable equilibria. Competition between the distinct solutions leads in these cases to the occurrence of a first-order phase transition, characterised by discontinuities in particle concentrations.
The previous analysis of Ref. [34], largely by numerical means, explored the general structure of the phase diagram in the interaction- and mass-balanced case with repulsive interactions. The major findings of that study are summarized as follows. For $D>0$, the phase diagram hosts, in addition to the normal phase, a phase involving condensate of only type-1 particles (BEC1), only type-2 particles (BEC2) and condensates of both species (BEC12), where all the involved phase transitions are continuous. This situation is schematically depicted in Fig. 1 (the right panel). In contrast, for $D<0$, there is no BEC12 phase, and the transition between BEC1 and BEC2 is of first order and is accompanied by phase separation; condensation of single species of particles may occur either via a first order or continuous transition depending on system parameters. In addition, for sufficiently large, positive interspecies couplings $a_{12}$, there occurs a line of liquid-gas-type phase transitions within the normal phase; for a schematic illustration, see Fig. 1 (left panel). This picture was primarily obtained through numerical analysis of Eq. (5,9,10) for situations with neither mass nor interaction imbalance and repulsive interparticle forces. The role of the sign of $D$ had been recognized in much earlier studies on this topic. Specifically, at $T=0$, varying $D$ across zero leads to a mixing-demixing transition in the system. For temperatures above zero with $D=0$, the transition between the BEC1 and BEC2 phases is expected to be of a multicritical nature [34].
We presently systematize, generalize, and extend that analysis to parameter ranges not addressed before, also providing a much more thorough analytical understanding. We begin by discussing the possibility of attractive couplings.
@@ III Attractive interspecies interactions
We first observe that the one-component counterpart of the system given by Eq. (1) is well-defined only for non-negative values of the interaction coupling. In regard to the presently studied two-component mixture, we note that the structure of the saddle-point equations [Eq. (9,10)] does not impose immediately obvious constraints on the values of the interaction couplings $\{a_{1},a_{2},a_{12}\}$. In fact, as we will demonstrate below, while $a_{1}$ and $a_{2}$ must remain positive to prevent collapse, the system can still be stable for sufficiently small negative values of $a_{12}$, as the repulsive intraspecies interactions can overcome the impact of the attractive interspecies coupling. To identify restrictions on the values of these parameters stemming from the system stability, we reexamine the structure of the integrals of Eq. (3,4). We first point out [49, 48, 34] that the result is independent of the arbitrary parameters $\alpha_{1}$ and $\alpha_{2}$ occurring in the integration limits. These parameters are taken to be negative to guarantee that the function $\Phi(t_{1},t_{2})$ can (at least for repulsive interactions) be analytically continued into the region of the complex plane encompassing the integration contour. To ensure that the Bose function $g_{5/2}(x)$ and the logarithmic function $\ln(x)$ are well-defined, both $|e^{\beta t_{1}}|$ and $|e^{\beta(\frac{a_{12}}{a_{1}}t_{1}+t_{2})}|$ must be less than one. This condition is satisfied provided that the real parts of $t_{1}$ and $(\frac{a_{12}}{a_{1}}t_{1}+t_{2})$ are negative. When $D>0$, it is straightforward to see that this can be guaranteed by appropriate choice of $\alpha_{1}$ and $\alpha_{2}$ (irrespective of the sign of $a_{12}$). On the other hand, when $D<0$ and $a_{12}<0$, it is impossible to fulfil the above conditions. To see this, we express $t_{1}$ as the sum of its real and imaginary parts, $t_{1}=a+ib$. Fulfillment of the inequality $|e^{\beta t_{1}}|<1$ requires that $a<0$. However, the condition $|e^{\beta(\frac{a_{12}}{a_{1}}t_{1}+t_{2})}|<1$ requires $a$ to be positive, as $a_{12}<0$ and $t_{2}$ is integrated over the entire real axis. This leads to a contradiction, rendering it impossible to satisfy both conditions simultaneously when $D<0$. In consequence, the integrals in Eq. (3, 4, 5) remain well defined also for attractive interspecies interactions but only as long as $D>0$. We also note that the intraspecies couplings $a_{1}$, $a_{2}$ must necessarily be positive.
@@
Figure 2: Projection of the phase diagram on the $(\mu_{1},\mu_{2})$ plane for a sequence of values of the interspecies coupling $a_{12}$ $\beta a_{12}\lambda_{1}^{-3}=0.9$ (blue dotted lines), $\beta a_{12}\lambda_{1}^{-3}=0.1$ (red dashed lines), and $\beta a_{12}\lambda^{-3}_{1}=-0.7$ (grey lines)] for sufficiently weak interspecies coupling $a_{12}$ such that $D>0$. In each of the plotted cases the general structure of the phase diagram is analogous to the one of the right diagram of Fig. 1; the BEC12 phase occupies the wedge-shaped region characterized by large values of $\mu_{1}$ and $\mu_{2}$ (corresponding also to high densities). The system hosts a quadruple point, where the BEC12, BEC1, BEC2 and the normal phase coexist. For $a_{12}>0$ and $D=0$ the BEC12 phase does not exist (the corresponding wedge angle $\phi=0$); the BEC12 wedge opens upon decreasing $a_{12}$, its angle reaches the value $\phi=90$°  for $a_{12}=0$ and approaches $\phi=180$°  in the limit $a_{12}\to-\sqrt{a_{1}a_{2}}^{+}$ (where $D\to 0^{+}$ and the system loses its stability). The plot parameters are $\beta a_{1}\lambda_{1}^{-3}=\beta a_{2}\lambda_{1}^{-3}=\lambda_{1}/\lambda_{2}=1$. All the phase transitions are continuous (see Sec. IV).
For $D>0$ and relatively small and positive values of $a_{12}$, the BEC12 phase is stable in the acute wedge-shaped region characterized by large values of $\mu_{1}$ and $\mu_{2}$ (which also correspond to high densities) - see Fig. 2. The system hosts a quadruple point, where the BEC12 phase coexists with the BEC1, BEC2, as well as with the normal phase. All the phase transitions involved here are continuous (see Sec. IV). Upon reducing $a_{12}$, the position of the quadruple point becomes shifted to lower values of $(\mu_{1},\mu_{2})$, and the BEC12 wedge progressively opens. The wedge angle $\phi$ reaches the value 90°  for $a_{12}=0$ and is further increased as $a_{12}$ becomes negative. The limiting value $a_{12}\to-\sqrt{a_{1}a_{2}}^{+}$ (where $D\to 0^{+}$) corresponds to the fully open wedge ($\phi\to$180°). This picture may be understood realizing that at the BEC2-BEC12 transition one has [see Eq. (9, 10)]:
@@
The analysis of attractive interspecies interactions concludes here; henceforth, we will consider $a_{12}>0$, where additional structures of the phase diagram appear.
@@ IV Tricritical points for condensation
In this section, we analyse the conditions for realising condensation as a first-order transition. Characterised by discontinuities of $n_{1}$, and $n_{n}$ as functions of $\mu_{1}$, $\mu_{2}$ and $T$. We demonstrate the impossibility of such a scenario for $D>0$ (a fact used in the analysis of the previous section). We also show that $D<0$ is a necessary but not sufficient condition for the appearance of first-order condensation in the system’s phase diagram. This scenario will, however, always be realised for $T$ sufficiently low.
We focus on the transition line between the normal and BEC2 phases; the normal-BEC1 case may be addressed in a fully analogous manner. For sufficiently large negative values of $\mu_{1}$ ($\beta\mu_{1}\gg 1$), the concentration of type-1 particles is negligibly low and properties of the mixture are equivalent to those of the one-component system. In particular, there is only one solution to the saddle-point equations; condensation of type-2 particles occurs via a continuous transition, and the corresponding phase boundary follows Eq. (14). It may, however, happen [34] that, as one moves along the line defined by Eq. (14), additional solutions to the saddle-point equation appear and Eq. (14) begins to represent an unstable equilibrium. This scenario requires an appearance of a tricritical point in the phase diagram ($P_{2,c}$ in Fig. 1), which is reflected by a change of the character of the extremum of $\Phi(n_{1}(n_{2}),n_{2})$ at the critical density. This, in turn, implies an existence of a point on the line given by Eq.(14) where the second derivative of $\Phi(n_{1}(n_{2}),n_{2})$ with respect to $n_{2}$ vanishes. Using Eq. (13), we view $\Phi(n_{1},n_{2})$, as a function of $n_{2}$ only and we introduce $\tilde{\Phi}(n_{2})\equiv\Phi(n_{1}(n_{2}),n_{2})$. The condition for vanishing of the second derivative of $\tilde{\Phi}(n_{2})$ then may be written as:
@@
We emphasize that the above discussion implies that the tricritical points may occur exclusively for negative values of the determinant $D$. However, a negative $D$ does not necessarily imply that Eqs. (20, 21) indeed represent a meaningful tricritical point on the phase diagram. The additional requirement is that the putative tricritical point of Eq. (20, 21) be located to the left (i.e. correspond to lower values of the chemical potentials) from the point of intersection of the putative second-order phase boundaries [given by Eq. (14) and an analogous equation for the normal-BEC1 transition]. Numerical analysis indicates that this condition becomes fulfilled for sufficiently low temperatures and/or $D$ sufficiently close to zero. This is easily proven in the balanced case, where $a_{1}=a_{2}=a$ and $\lambda_{1}=\lambda_{2}=\lambda$. In this case, the condition for Eq. (20, 21) to represent a physical tricritical point reduces to $\mu_{1,tri}<\mu_{2,tri}$. After applying Eqs. (20, 21) and recalling that $n_{2,c}=\frac{1}{\lambda_{2}^{3}}\zeta\left(3/2\right)$, this condition may be recast as: